OpenStudy (cookie_2046):

What is the value of expression? 6^−7 • 6^4 A. 6−3 B. 6−11 C. 6−28 D. 611

OpenStudy (astrophysics):

\[\huge 6^{(-7)} \times 6^4\]

OpenStudy (cookie_2046):

yes that.

OpenStudy (astrophysics):

Use this, \[\huge x^m \times x^n = x^{m+n}\]

OpenStudy (astrophysics):

When you multiply and have exponents you add the exponents and if you're dividing you subtract the exponents.

OpenStudy (♪chibiterasu):

When you multiply expressions with exponents, simply add the exponents and multiply the base value.

OpenStudy (♪chibiterasu):

Which is just a description of the formula @Astrophysics shared with you.

OpenStudy (astrophysics):

\[\huge 6^{(-7)} \times 6^4 \implies \frac{ 6^4 }{ 6^7 }\] I don't know if that will make it easier for you

OpenStudy (cookie_2046):

I'm just confused, i don't know if i subtract the exponent or add it

OpenStudy (astrophysics):

Notice I just used \[\huge x^{-n} = \frac{ 1 }{ x^n }\] and it doesn't matter, all you're really doing is -7+4

OpenStudy (♪chibiterasu):

I jsut told you to add them lol.

OpenStudy (kidrah69):

Well its multiplacation so you would add

OpenStudy (kidrah69):

When its division you subtract

OpenStudy (kidrah69):

No -7 + 4

OpenStudy (kidrah69):

Remeber its a -7

OpenStudy (astrophysics):

How did you get 11?

OpenStudy (kidrah69):

Not a postive 7

OpenStudy (kidrah69):

No whats -7+4?

OpenStudy (kidrah69):

add 4 to NEGATIVE 7

OpenStudy (astrophysics):

How are you getting b can you please show what you're doing, if we just give you the answer that will not help you, as kidrah is asking you what's -7+4 this is simple arithmetic.

OpenStudy (cookie_2046):

it would be 0

OpenStudy (kidrah69):

No

OpenStudy (kidrah69):

Seems like you are trolling if you dont get this use a calculator

OpenStudy (astrophysics):

\[\huge -7 + 4 = ???\]

OpenStudy (cookie_2046):

-3 i mean

OpenStudy (astrophysics):

Yeah. So what will be your answer then

OpenStudy (kidrah69):

\(\huge\color{red}{CONGRATS!!}\)

OpenStudy (kidrah69):

Good job

OpenStudy (cookie_2046):

a.

OpenStudy (kidrah69):

yep

OpenStudy (cookie_2046):

i'm bad in math sorry

OpenStudy (kidrah69):

its okay but next time try a calculator :)

OpenStudy (astrophysics):

It's ok, and no do not use a calculator, I suggest you do some arithmetic problems instead that will benefit you rather than a silly calculator :P

OpenStudy (kidrah69):

Well for problems like this if he doesnt know then a calculator would be helpful :P

OpenStudy (kidrah69):

But yeah they should learn without one too :P yr right :3

OpenStudy (♪chibiterasu):

This post is funny.